OpenStudy (anonymous):

!!Urgent help needed!! Find an explicit solution of the given initial-value problem. (x^2)(dy/dx) = y-xy, y(-1) = -1

OpenStudy (anonymous):

@e.mccormick Do you think you could help me really quick?

OpenStudy (anonymous):

I have the answer, Im just lost in the steps. A quick walkthrough is all i need.

ganeshie8 (ganeshie8):

you may solve this by 'separating variables'

ganeshie8 (ganeshie8):

(x^2)(dy/dx) = y-xy (x^2)(dy/dx) = y(1-x)

OpenStudy (anonymous):

I got that part, i dont know how to solve for c

ganeshie8 (ganeshie8):

solving c is easy, you just need to plugin the point (-1, -1) in general solution. wats the general solution u got ?

OpenStudy (anonymous):

C=-e^(-2)

ganeshie8 (ganeshie8):

im asking about general solution..

OpenStudy (anonymous):

thats where im stuck. I got y=ce^(-1/x-x)

ganeshie8 (ganeshie8):

dx(1-x)/x^2 = dy/y x^-2 - 1/x dx = 1/y dy -1/x - ln |x| = ln |y| + C ln |xy| = -1/x + c xy = e^(-1/x+c) y = e^(-1/x+c)/x

ganeshie8 (ganeshie8):

^^ i am getting that, next rewrite constant e^c as A

ganeshie8 (ganeshie8):

y = e^(-1/x+c)/x y = Ae^(-1/x)/x plugin the point (-1, -1) and solve A

ganeshie8 (ganeshie8):

y = Ae^(-1/x)/x -1 = Ae^(-1/-1) / -1 1 = Ae A = 1/e

OpenStudy (anonymous):

book says: y = (e^-(1+1/x))/x

ganeshie8 (ganeshie8):

yes thats correct

ganeshie8 (ganeshie8):

y = Ae^(-1/x)/x plugin A = 1/e and simplify